a planet of mass 1/10 of the earth and radius of 1/3 of the earth. If a person can throw a stone on earth surface to a height of 90m, he will able to throw the stone on that planet to a height

Grade:11

2 Answers

Vikas TU
7 years ago
g constant on earth that is,
g = GM/R
 g constat at the planet that is,
g’ = 3GM/10R
relation b/w them is:
g’ = 3g/10
Now energy would remain contant.
Hence,
mg*90 = m*3g*h/10
we get,
h =  300 m.
Namisha
5 years ago
We have;
g=Gm/R^2
'g' on the other planet using the above formula is
g=Gm/10(R/3)^2
   =9Gm/10R^2
Which is...g of other planet is equals to 9/10 times of that of the g of Earth.
So; as the energy would remain constant
Use mgh..
m*g*90=m*9g*h/10
So from here h=100m
 
So h=100m is the height at which we will be able to throw the stone on that planet 
:-)